A firm is considering three possible one-year investments, which we will name X, Y, and Z.

     • Investment X would cost $10 million now and would return $11 million next year, for a net gain of $1 million.

     • Investment Y would cost $100 million now and would return $105 million next year, for a net gain of $5 million.

     • Investment Z would cost $1 million now and would return $1.2 million next year, for a net gain of $200,000.

The firm currently has $150 million of cash on hand that it can loan out at 15 percent interest. Which of the three possible investments should it undertake?

  • X only

  • Y only

  • Z only

  • X and Y

  • X and Z

  • X, Y, and Z

  • None
Answer #1

Answer:

Correct answer is:

Z only

Explanation:

Opportunity cost of capital is =15%

NPV of investment X = 11000000 / (1 + 15%) - 10000000 = -$434,782.61

NPV of investment Y = 105,000,000 / (1 + 15%) - 100,000,000 = - $8,695,652.17

NPV of investment Z = 1200000 / (1 + 15%) - 1000000 = $43478.26

We observe from NPVs of 3 investment options that only investment Z gives positive NPV and hence the firm should take up only investment Z.

Hence option C is correct and all other options A, B, D, E, F and G are incorrect.